# Cross-Lingual Unlearning of Selective Knowledge in Multilingual Language Models

### v1

@@ Abstract
Minseok Choi       Kyunghyun Min       Jaegul Choo KAIST AI     Ajou University {minseok.choi,jchoo}@kaist.ac.kr khmin1104@ajou.ac.kr
@@ 1 Introduction
Privacy regulations such as the Right to be Forgotten (RTBF) Rosen (2011), the European Union’s General Data Protection Regulation (GDPR) Hoofnagle et al. (2019), and the United States’ California Consumer Privacy Act (CCPA) Pardau (2018) mandate that individuals have the right to request the deletion of their data from databases, which extends to data held within machine learning (ML) models. Additionally, the Writers Guild of America strike in 2023 highlighted increasing concerns regarding the copyright content generated by large language models (LLMs) WGA (2023). To comply with such issues, significant attention has been directed towards the task of machine unlearning (MU), which involves removing the influence of specific data points from ML models Cao and Yang (2015). Despite the critical necessity of the task, mitigating the influence of data samples on billions of model parameters presents an immense challenge. The most definitive method is exact unlearning, which necessitates retraining ML models entirely from scratch, utilizing the residual training dataset after excising the specified data points. However, this method is computationally prohibitive and not feasible, particularly for LLMs. Therefore, the advancement of rapid approximate unlearning methodologies has emerged as a primary focus of contemporary research efforts.
@@
2 Related Work
2.1 Machine Unlearning
@@ 2.2 Knowledge Unlearning
Jang et al. (2023) introduced knowledge unlearning, aimed at preventing language models from generating specific token sequences. They proposed a straightforward method by inverting the original training objective of minimizing the negative log-likelihood of the token sequences. To maintain the performance of the remaining knowledge, Wang et al. (2023) and Chen and Yang (2023) employed the Kullback-Leibler (KL) divergence loss, minimizing the distributional differences between the original and unlearned models on the retained data. Our approach builds on these methods but differs in its focus. While the previous works targeted monolingual models like DistilBERT Sanh et al. (2019) and T5 Raffel et al. (2020), we extend the unlearning process to a multilingual context.
2.3 Cross-Lingual Transfer
Multilingual pretrained language models Devlin et al. (2019); Conneau et al. (2020); Xue et al. (2021); Lin et al. (2022); Le Scao et al. (2023) have shown to exhibit remarkable cross-lingual transfer across various tasks by leveraging shared semantic spaces and joint training techniques to bridge language gaps. However, Xu et al. (2023) demonstrated that editing knowledge in one language does not propagate to others and thus introduced cross-lingual model editing, a technique using random sampling of languages to improve model adaptability and robustness in a multilingual context. Additionally, Qi et al. (2023) investigated the cross-lingual consistency of factual knowledge in multilingual models, finding that factual knowledge does not remain consistent across languages, but only when languages share a larger portion of vocabulary. Building on these advancements, our work employs multilingual language models to investigate the cross-lingual transfer of machine unlearning and proposes an effective method to unlearn specific information across languages, addressing the need for precise and reliable information removal in a multilingual context.
3 Methodology
3.1 Problem Definition
Given a token sequence $\mathbf{x}=\{x\}_{i=1}^{T}$ in the training dataset $\mathcal{D}=\{\mathbf{x}\}_{i=1}^{N}$, the task of knowledge unlearning is to safely remove the influence of a subset of data $\mathcal{D}_{f}$ from a trained machine learning model such that the model behaves as if the removed data had never been part of the training process, thus maintaining the model performance for the rest of the dataset. Conventionally, the data to be forgotten $\mathcal{D}_{f}$ is expressed as the forget set, while the data to be retained $\mathcal{D}_{r}$ is named as the retain set. For simplicity, we consider the standard case where $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ represent the whole training dataset and are mutually exclusive; that is, $\mathcal{D}_{f}\cup\mathcal{D}_{r}=\mathcal{D}$ and $\mathcal{D}_{f}\cap\mathcal{D}_{r}=\varnothing$. The objective is to adjust the model parameters $\theta$ such that the updated parameters $\theta^{\prime}=S(\theta;\mathcal{D}_{f})$ reflect the removal of $\mathcal{D}_{f}$. This unlearning (scrubbing) function $S$ ensures the model behaves as if trained solely on $\mathcal{D}_{r}$, effectively forgetting $\mathcal{D}_{f}$ while maintaining performance on $\mathcal{D}_{r}$. Extending to a multilingual context, the definition must hold for all datasets across languages $\mathcal{Z}=\{z\}_{i=1}^{Z}$.
3.2 Knowledge Unlearning
The primary objective in language modeling is to minimize the log-likelihood of token sequences, training the model to predict the next word in a sequence accurately. Knowledge unlearning Jang et al. (2023) involves negating this objective to remove specific learned information from the model. Instead of reinforcing certain sequences, unlearning aims to decrease their probability by maximizing their log-likelihood:
$\mathcal{L}_{f}=-\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(x_{t}|x_{<t}),$ (1)
@@ 3.3 Language-Adaptive Unlearning
After forgetting a subset of data, many previous works highlight the critical need to retain the rest of the knowledge explicitly Wang et al. (2023); Chen and Yang (2023). This involves adjusting the model so that its performance on retained data aligns closely with the original model as if the forgotten samples never existed. Formally, this can be expressed as minimizing the KL divergence between the original model and the unlearned model on the retained data:
$\mathcal{L}_{LT}=\frac{1}{T}\sum_{t=1}^{T}D_{\text{KL}}(p_{\theta_{0}}(\cdot|x_{<t})\;\|\;p_{\theta}(\cdot|x_{<t})),$ (2)
@@
$\mathcal{L}_{LM}=\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(x_{t}|x_{<t}).$ (3)
This represents the positive counterpart of Equation 1. Employing an adaptive weighting scheme, we can combine the language teaching loss and the language modeling loss:
$\mathcal{L}_{r}=\kappa\cdot\mathcal{L}_{LT}+(1-\kappa)\cdot\mathcal{L}_{LM},$ (4)
where $\kappa=\frac{1}{T}\sum_{t=1}^{T}p_{\theta_{0}}(\cdot|x_{<t})$ represents the confidence of the teacher in token sequence $\mathbf{x}$ for the given language. This implies that the student learns from the teacher when the teacher’s confidence is high; otherwise, the student learns independently to retain examples. Finally, combining all losses, we obtain
@@
where $\lambda$ is a scaling hyperparameter. In practice, we follow Kurmanji et al. (2023), alternating the updates for the forget set and the retain set to optimize min-max terms in $\mathcal{L}$ more stably. Furthermore, this objective supports token-level unlearning, indicating that cross-lingual transfer will only occur if the languages share the same vocabulary, as noted by Qi et al. (2023). To facilitate fast and effective cross-lingual unlearning, we randomly sample languages for token sequences in both the forget and retain sets, following Xu et al. (2023). We demonstrate in Section 5 that this approach achieves comparable performance to unlearning one language at a time, with significantly improved efficiency.
@@ 4.2 Baselines
We compare our framework with several strong unlearning approaches and various baselines: Original: The “original” model without any unlearning applied. GradAscent+: This method begins with the original model and finetunes it on both the retain and forget sets, using gradient ascent on the latter. Previous work Jang et al. (2023) examined a weaker baseline that only trains on the forget set with gradient ascent. We enhance GradAscent+ to achieve a better balance between retention and forgetting. NegTaskVector+: This approach also starts from the original model but finetunes two separate models, one on the forget set and another on the retain set. During inference, the weights of the forget-set-tuned model are negated, while the retained weights are added. Prior research Ilharco et al. (2023) explored a weaker baseline training only on the forget set. Our refined version includes explicit retention tuning. Oracle: Serves as a reference point where our proposed method is applied one language at a time. This represents the “pseudo” upper bound performance of our approach, achieved inefficiently as the number of languages increases, i.e., $O(Z)$. We do not directly compare with other teacher-student frameworks for unlearning Chundawat et al. (2023); Kurmanji et al. (2023), as their training objectives involve a classification loss to forget a class label. Instead, we evaluate our adaptive unlearning scheme against the general knowledge distillation framework to demonstrate its effectiveness, as detailed in Section 5.3.
@@ 6 Conclusion
Despite the robust findings presented in our paper, certain limitations warrant discussion. The datasets used to explore multilingual unlearning in this study, namely FLORES and BMLAMA, are in the general domain. This is due to the scarcity of multilingual parallel datasets, especially within specific domains such as privacy data. This challenge mirrors those seen in computer vision, where datasets like CIFAR and MNIST, although unrelated to privacy, are used due to the difficulty in obtaining privacy-specific data. Future research should focus on inventing and benchmarking real or synthetic privacy data in multilingual settings to address these gaps. Additionally, our research was constrained by GPU resources, preventing us from testing models with 7B parameters or more. Investigating whether our conclusions hold for larger-scale models is a promising avenue for future work.
@@ Appendix A Additional Details for LingTea
A.1 Hyperparameters
@@ B.1 Token Sequence Unlearning Results for Each Language
We report the per-language performance of unlearning token sequences in FLORES-200 across compared models in Tables 6, 7, 8, and 9.
@@ B.2 Factual Knowledge Unlearning Results for Each Language
We report the per-language performance of unlearning factual knowledge in BMLAMA-53 across compared models in Tables 10, 11, 12, and 13.

### v2

@@ Abstract
Minseok Choi     Kyunghyun Min     Jaegul Choo KAIST AI    Yonsei University {minseok.choi,jchoo}@kaist.ac.kr khmin1104@yonsei.ac.kr
@@ 1 Introduction
Privacy regulations such as the Right to be Forgotten (RTBF) Rosen (2011), the European Union’s General Data Protection Regulation (GDPR) Hoofnagle et al. (2019), and the United States’ California Consumer Privacy Act (CCPA) Pardau (2018) mandate that individuals have the right to request the deletion of their data from databases, which extends to data held within machine learning (ML) models. Additionally, the Writers Guild of America strike in 2023 highlighted increasing concerns regarding the copyrighted content generated by large language models (LLMs) WGA (2023). To comply with such issues, significant attention has been directed towards the task of machine unlearning (MU), which involves removing the influence of specific data points from ML models Cao and Yang (2015). Despite the critical necessity of the task, mitigating the influence of data samples on billions of model parameters presents an immense challenge. The most definitive method is exact unlearning, which necessitates retraining ML models entirely from scratch, utilizing the residual training dataset after excising the specified data points. However, this method is computationally prohibitive and not feasible, particularly for LLMs. Therefore, the advancement of rapid approximate unlearning methodologies has emerged as a primary focus of contemporary research efforts.
@@
2 Problem Definition
Knowledge Unlearning
Given a token sequence $\mathbf{x}=\{x\}_{i=1}^{T}$ in the training dataset $\mathcal{D}=\{\mathbf{x}\}_{i=1}^{N}$, the task of knowledge unlearning is to safely remove the influence of a subset of data $\mathcal{D}_{f}$ from a trained machine learning model such that the model behaves as if the removed data had never been part of the training process, thus maintaining the model performance for the rest of the dataset. Conventionally, the data to be forgotten $\mathcal{D}_{f}$ is expressed as the forget set, while the data to be retained $\mathcal{D}_{r}$ is named as the retain set. For simplicity, we consider the standard case where $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ represent the whole training dataset and are mutually exclusive; that is, $\mathcal{D}_{f}\cup\mathcal{D}_{r}=\mathcal{D}$ and $\mathcal{D}_{f}\cap\mathcal{D}_{r}=\varnothing$. The objective is to adjust the model parameters $\theta$ such that the updated parameters $\theta^{\prime}=S(\theta;\mathcal{D}_{f})$ reflect the removal of $\mathcal{D}_{f}$. This unlearning (scrubbing) function $S$ ensures the model behaves as if trained solely on $\mathcal{D}_{r}$, effectively forgetting $\mathcal{D}_{f}$ while maintaining performance on $\mathcal{D}_{r}$.
Multilingual Unlearning
Extending to a multilingual context, the above definition must hold for the dataset $\mathcal{D}$ across all possible languages. While ideally, unlearning should occur across all existing languages, our experiments focus on a predefined set of languages $\mathcal{Z}=\{z\}_{i=1}^{Z}$ for feasibility. Consequently, we assume a parallel dataset with forget sets $\mathcal{D}_{f}^{1},\mathcal{D}_{f}^{2},...,\mathcal{D}_{f}^{Z}]$ and retain sets $\mathcal{D}_{r}^{1},\mathcal{D}_{r}^{2},...,\mathcal{D}_{r}^{Z}]$. We define successful multilingual unlearning as the effective forgetting of parallel samples across all forget sets while retaining parallel samples across all retain sets.
3 Methodology
@@ 3.1 Knowledge Unlearning
The primary objective in language modeling is to minimize the negative log-likelihood of token sequences, training the model to predict the next word in a sequence accurately. Knowledge unlearning Jang et al. (2023) involves negating this objective to remove specific learned information from the model. Instead of reinforcing certain sequences, unlearning aims to decrease their probabilities by maximizing their negative log-likelihood, which can be understood as equivalent to removing the negative sign:
$\mathcal{L}_{f}=\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(x_{t}|x_{<t}),$ (1)
@@ 3.2 Language-Adaptive Unlearning
After forgetting a subset of data, many previous works highlight the critical need to retain the rest of the knowledge explicitly Wang et al. (2023); Chen and Yang (2023); Lee et al. (2024). This involves adjusting the model so that its performance on retained data aligns closely with the original model as if the forgotten samples never existed. Formally, this can be expressed as minimizing the KL divergence between the original model and the unlearned model on the retained data:
$\mathcal{L}_{\text{LT}}=\frac{1}{T}\sum_{t=1}^{T}D_{\text{KL}}(p_{\theta_{0}}(\cdot|x_{<t})\;\|\;p_{\theta}(\cdot|x_{<t})),$ (2)
@@
$\mathcal{L}_{\text{LM}}=-\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(x_{t}|x_{<t}).$ (3)
This represents the negative counterpart of Equation 1. Employing an adaptive weighting scheme, we can combine the language teaching (LT) loss and the language modeling (LM) loss:
$\mathcal{L}_{r}=\kappa\cdot\mathcal{L}_{\text{LT}}+(1-\kappa)\cdot\mathcal{L}_{\text{LM}},$ (4)
where $\kappa=\frac{1}{T}\sum_{t=1}^{T}p_{\theta_{0}}(\cdot|x_{<t})$ represents the confidence of the teacher in token sequence $\mathbf{x}$ for the given language. This implies that the student learns from the teacher when the teacher’s confidence is high; otherwise, the student learns independently to retain examples. One could argue that this language-adaptive scheme may be limited by simply doing more pretraining on low-resource languages. However, we contend that while additional pretraining might enhance the model’s overall confidence in tokens of that particular language, it could still struggle to capture rare tokens, whether in high- or low-resource languages. The efficacy of our approach mainly comes from its ability to adjust weights for all tokens, whether they are frequent or rare and whether they are in high- or low-resource languages, enabling more precise weighting and fine-grained control over individual tokens.
3.3 Training
Combining the forgetting and retaining losses, we obtain the following:
@@ 3.3 Training
where $\lambda$ is a scaling hyperparameter. In practice, we follow Kurmanji et al. (2023), alternating the updates for the forget set and the retain set to optimize min-max terms in $\mathcal{L}$ more stably. Furthermore, to facilitate fast and effective multilingual unlearning, we randomly sample languages for token sequences in both the forget and retain sets, following Xu et al. (2023). Specifically, we employ $\mathcal{D}_{f}^{z}$ and $\mathcal{D}_{r}^{z}$ where $z$ is randomly sampled from $\mathcal{Z}$ during training. We demonstrate in §5 that this approach achieves comparable performance to unlearning one language at a time, with significantly improved efficiency.
@@ 4.2 Baselines
We compare our framework with several strong unlearning approaches and various baselines: Original: The “original” model without any unlearning applied. GradAscent+: This method begins with the original model and finetunes it on both the retain and forget sets, using gradient ascent on the latter. Previous work Jang et al. (2023) examined a weaker baseline that only trains on the forget set with gradient ascent. We enhance GradAscent+ to achieve a better balance between retention and forgetting. NegTaskVector+: This approach also starts from the original model but finetunes two separate models, one on the forget set and another on the retain set. During inference, the weights of the forget-set-tuned model are negated, while the retained weights are added. Prior research Ilharco et al. (2023) explored a weaker baseline training only on the forget set. Our refined version includes explicit retention tuning. Oracle: Serves as a reference point where our proposed method is applied one language at a time. This represents the “pseudo” upper bound performance of our approach, achieved inefficiently as the number of languages increases, i.e., $O(Z)$. We do not directly compare with other teacher-student frameworks for unlearning Chundawat et al. (2023); Kurmanji et al. (2023), as their training objectives involve a classification loss to forget a class label. Instead, we evaluate our adaptive unlearning scheme against the general knowledge distillation framework to demonstrate its effectiveness, as detailed in §5.3.
4.3 Evaluation Metrics
@@ 5.5 Scaling the Number of Samples to Forget
6 Related Work
6.1 Machine Unlearning
6.2 Knowledge Unlearning
Jang et al. (2023) introduced knowledge unlearning, aimed at preventing language models from generating specific token sequences. They proposed a straightforward method by inverting the original training objective of minimizing the negative log-likelihood of the token sequences. To maintain the performance of the remaining knowledge, various works Wang et al. (2023); Chen and Yang (2023); Lee et al. (2024) employed the Kullback-Leibler (KL) divergence loss, minimizing the distributional differences between the original and unlearned models on the retained data. Our approach builds on these methods but differs in its focus. While the previous works targeted monolingual models like DistilBERT Sanh et al. (2019) and T5 Raffel et al. (2020), we extend the unlearning process to a multilingual context.
6.3 Cross-Lingual Transfer
Multilingual pretrained language models Devlin et al. (2019); Conneau et al. (2020); Xue et al. (2021); Lin et al. (2022); Le Scao et al. (2023) have shown to exhibit remarkable cross-lingual transfer across various tasks by leveraging shared semantic spaces and joint training techniques to bridge language gaps. However, Xu et al. (2023) demonstrated that editing knowledge in one language does not propagate to others and thus introduced cross-lingual model editing, a technique using random sampling of languages to improve model adaptability and robustness in a multilingual context. Additionally, Qi et al. (2023) investigated the cross-lingual consistency of factual knowledge in multilingual models, finding that factual knowledge does not remain consistent across languages, but only when languages share a larger portion of vocabulary. Building on these advancements, our work employs multilingual language models to investigate the cross-lingual transfer of machine unlearning, revealing that current unlearning methods lack this capability. In response, we propose an effective approach for unlearning specific information across languages, addressing the need for precise and reliable information removal in a multilingual context.
@@ 7 Conclusion
Despite the robust findings presented in our paper, certain limitations warrant discussion. The datasets used to explore multilingual unlearning in this study, namely FLORES and BMLAMA, are in the general domain. This is due to the scarcity of multilingual parallel datasets, especially within specific domains such as privacy data. This challenge mirrors those seen in computer vision, where datasets like CIFAR and MNIST, although unrelated to privacy, are used due to the difficulty in obtaining privacy-specific data. Future research should focus on inventing and benchmarking real or synthetic privacy data in multilingual settings to address these gaps. Additionally, our research was constrained by GPU resources, preventing us from testing models with 7B parameters or more. Investigating whether our conclusions hold for larger-scale models is a promising avenue for future work. Lastly, our work does not handle indirect cases; that is, our framework assumes a clear data separation between forget and retain sets. Before addressing this issue, we must consider: 1) whether we aim to remove the target tokens/sequences present in other texts, and 2) whether we should also unlearn indirect expressions. In both cases, data augmentation could be a potential remedy in enhancing the model’s robustness against indirect cases, though it may introduce unforeseen consequences. We hope future work can address more details.
Acknowledgements
We sincerely thank ChaeHun Park for sharing his valuable knowledge during our discussions. Additionally, we are grateful to the anonymous reviewers for their insightful feedback. This work was supported by the Institute for Information & Communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program (KAIST)), the artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City, and Samsung Electronics Co., Ltd.
@@ Appendix A Additional Details for LingTea
  Forget Set Test Set
en high-src low-src en high-src low-src
XGLM-564M GA+KL 23.0 7168.7 24.1 2252.5 21.2 2156.0 30.8 1192.7 29.8 618.8 25.1 853.4
XGLM-2.9B GA+KL 22.6 1583765.3 26.9 1875148.9 26.5 269117.1 37.2 26177.9 37.5 12710.0 33.8 26767.7
BLOOM-560M GA+KL 20.2 171659.1 21.0 28560.2 15.1 318697.5 28.5 26366.8 27.8 5924.2 18.5 46637.4
BLOOM-3B GA+KL 21.3 359746.9 21.5 138652.5 17.5 78245.9 35.5 334.8 34.6 206.9 25.3 785.1
Table 4: Performance of unlearning multilingual token sequences on FLORES-200, comparing with GA+KL.
A.1 Additional Baseline Results
We also report the GA+KL baseline, which unlearns the forget set using gradient ascent (GA) and maintains the retaining performance using KL divergence. In the main experiment, we compare our framework with GradAscent+, which retains using gradient descent. However, since our framework is an adaptive mixture of the two, it is equally important to compare with the pure KL baseline. As shown in Table 4, we observe that our method LingTea consistently outperforms the GA+KL baseline, demonstrating the effectiveness of our approach. This exhibits that it is very crucial to adaptively obey the teacher when unlearning in a multilingual context.
A.2 Hyperparameters
@@ B.1 Token Sequence Unlearning Results for Each Language
We report the per-language performance of unlearning token sequences in FLORES-200 across compared models in Tables 7, 8, 9, and 10.
@@ B.2 Factual Knowledge Unlearning Results for Each Language
We report the per-language performance of unlearning factual knowledge in BMLAMA-53 across compared models in Tables 11, 12, 13, and 14.